• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kai

    Log entry time 13:10:08 on October 05, 2009

    Entry number 293276

    keyword=daily WAC report

    1. The run been analyzed is from #12933~#12967.

    2. Among them, #12934~#12943, moller test, pedstal run. #12948~#12953 no beam, LED linearity test. All of these set as non production.

    3. #12960~#12965, right arm dipole was down and took time ramping up. #12963,64,65 magnet is oscillating, the left arm data should be OK. However, I don't know how to get this out, so I also set these runs as non-production.

    Total OK pairs 507.8K. Switch WAC work to Megan Friend.

    A copy of this log entry has been emailed to: paschke,mercado,mcnulty,pkpan,silwal,zhaobo,rom,dalton,mfriend